Given the following inverse demand and supply functions Supply p 40+30 Demand p 89-20 Solve for the equilibrium quantity - IIl units. (round your answer to two decimat places)
Answer #1
At equilibrium,
Supply = Demand
40 + 3q = 89 - 2q
5q = 89 - 40
5q = 49
q = 9.80
The equilibrium quantity is 9.80 units.